# Exercising the CCPA Opt-out Right on Android: Legally Mandated but Practically Challenging

### v5

@@ 6.2.2. Connecting to Ad Tracking Domains
Figure 9. Counts of apps in our app dataset connecting to the 29 most prevalent ad tracking domains that are subject to the CCPA opt-out right (Figure 5). There are no substantial differences between conditions.:
Evaluating the counts of apps in our app dataset that make ad tracking domain connections, we do not find substantial differences between the four conditions (Figure 9). However, there are minor differences. Notably, the count of apps that connect to the app-measurement.com domain in the AdID condition is higher (1,047 apps) than in the No AdID + GPC condition (808 apps). The reason for this difference could be a compounding effect where disabling the AdID and sending GPC signals each leads to a slight decrease. Overall, however, the number of apps connecting to ad tracking domains in the different conditions does not indicate that sending GPC signals is an effective means for exercising the CCPA opt-out right. Disabling the AdID did not change this result.
One other point is noteworthy. In our app dataset, all apps that connected to first-party servers also connected to third-party servers. However, 12% of apps connected exclusively to third-party servers. Those apps will receive GPC headers directly from a user’s device without reliance on redirects from a first-party server. For those apps we did not observe substantially different levels of GPC compliance. Thus, it does not appear that the communication model—receiving GPC signals via redirects from first parties vs direct GPC signals from a user’s device—plays a role for GPC compliance. This is in line with the GPC specification, which does not differentiate between the two (Zimmeck et al., 2025). Third parties are obligated to listen for GPC signals regardless from whom they receive them.
@@ 6.2.3. Disclosing Device Identifiers
Overall, the rates of device identifier disclosures under the four conditions are similar. The Android ID, serial number, and MAC address of our test device was not disclosed in any condition. In some requests to googleapis.com, we find our test device’s IMEI, which, however, is not correlated to any particular condition. Further, we find the public IP address of our test device in some third-party request or response bodies, though, again, not correlated to any particular condition. Interestingly, a small number of domains also still received the AdID in the No AdID and No AdID + GPC conditions. [^27] Three of the domains are Google domains — app-measurement.com, googleapis.com, and google-analytics.com — and three are non-Google domains — omtrdc.net, gadsme.com, and wizz.chat. We re-ran our analysis, which confirmed our results.
Since we quit each app and cleared its cache and data in between conditions, the AdID from a previous condition run could not have been left stored in the app’s storage at the time of the next condition run. Generally, the app could also not have written the AdID to other apps’ external storage on our test device. While our controller granted each app all permissions it requested at install time, the app would generally need direct user consent to read and write to other apps’ files (Android Open Source Project, 2024b). A plausible explanation is that the AdID was stored at the backend and associated with another identifier that was then synced once the app was run again in the new condition. We inquired with the respective companies but did not hear back.
@@ 7. Discussion: Implementing Opt-out Rights on Android and Improving User Privacy
Our results suggest that many Android apps do not provide California residents with an effective way to exercise their CCPA opt-out right. We see these shortcomings as a platform-level challenge. To broadly improve the privacy of Android users the Android platform would benefit from an opt-out setting to exercise opt-out rights under applicable law in place of its current AdID setting.
@@ 7.1. Android’s Opt-out Right Gap
The long-established Android AdID setting has no substantial effect on apps’ opt-out behavior (Section 6.2). Apps’ sharing and selling of personal information also generally continues despite sending GPC signals. At the app-level, our results suggest that many apps do not have a dedicated CCPA opt-out setting, and some that do are not propagating California residents’ opt-out preferences to third parties (Section 5.2). Given that 63%–82% of apps in our app dataset are required to respect the CCPA opt-out right (Section 6.1), our results suggest that many apps violate the CCPA as they do not enable California residents to exercise their opt-out right. It is generally difficult and in many case not possible for California residents to exercise their opt-out right.
There could be various reasons for why that is the case. First, traditionally the core business model of the ad online ecosystem has been directly opposed to respecting opt-out preferences (Zuboff, 2019). Until the recent enforcement actions by the Office of the California Attorney General (State of California Department of Justice, 2022b, a, 2025) the risk of being caught for not doing so was low. Some app operators may simply accept the risk. After all, dark patterns have been presumably proliferating because companies’ A-B testing has revealed them to be profit maximizing (Luguri and Strahilevitz, 2021). The same could be true for not providing opt-out functionality.
In our view, technical and ecosystem complexities also play a major role for the lack of CCPA opt-out right compliance. Opting out is a complex coordination problem with distributed responsibilities among first parties, ad networks, CMPs, and other parties, all of which are not aligned (Hils et al., 2021). These different responsibilities become difficult to manage without addressing them at the platform-level. Thus, in our view, our results indicate that app-level opt-out preferences are just not feasible. We need a platform-level opt-out right architecture.
@@ 7.2. Two Settings for One Task
As it stands, Android currently has two types of opt-out settings: the platform-level AdID and the app-level CCPA opt-out settings. The AdID setting is not intended for exercising the opt-out right per the CCPA or other privacy laws. However, according to Google, if a user has disabled the AdID, operators “may not use the advertising identifier for creating user profiles for advertising purposes or for targeting users with personalized advertising” (Google, 2024c). Similarly, the CCPA prohibits the sharing of personal information for purposes of cross-context behavioral advertising, defined as the “targeting of advertising to a consumer based on the consumer’s personal information” and which covers “consumer profiles” per the CCPA Regulations. [^28] CCPA §1798.140(k), CCPA Regulations, §7025(c)(1). Thus, the intended effect of the AdID setting is similar to what the CCPA opt-out right is intended to accomplish. However, despite this similarity the AdID setting is without legal meaning outside the contractual relationship that Google has with the app operators.
This division of the AdID setting from the opt-out right is based on its historical development predating the introduction of the CCPA (Section 2.2). In our view, it would benefit the Android platform if both were united and the opt-out right would be available on Android as a native setting. There should be only one opt-out setting on Android. The current division is confusing from a user perspective. Users have to both disable the AdID and perform app-level opt-outs to accomplish one task. If there were only one setting, it would give users a consistent experience and also relieve app operators from compliance burdens they may not be sufficiently equipped to tackle. Thus, we suggest to re-purpose the existing Android AdID setting as an opt-out right setting.
@@ 7.3. Re-purposing the AdID Setting
Disabling the AdID should be interpreted as exercising the opt-out right. This would immediately make the opt-out right on Android available to a wide audience in line with previous results suggesting that many users have interest in such setting (Zimmeck et al., 2023). Such re-purposing would unify and streamline the current two types of settings—the platform-level AdID setting and the app-level opt-out right settings—into one and lead to greater legal clarity. The current AdID API can be modified to return an opt-out signal (akin to GPC’s Sec-GPC: 1 header) instead of a string of zeros when a user opts out (Google, 2024d). Notably, opting out via GPC and the AdID setting are both intended as universal settings and have nearly identical scope (Appendix 10.2). However, ultimately, whether this uniform setting is a GPC setting or other opt-out right setting is secondary.
@@
Evolving the AdID setting and APIs towards the opt-out right requires a combination of technological and policy changes on Android. Creating a platform-level API that apps and their integrated libraries can query for the user’s opt-out status is a standard feature for an operating system or other platform-level software layer. To make users aware of their opt-out right, Android could surface a prompt for universally opting out as part of the setup of a new or reset Android device. Android could also surface such prompt upon the first start of an app that is selling or sharing personal information. Such prompt should contain an option for applying a user’s preference universally across all apps, e.g., via generalizable active privacy choice (Zimmeck et al., 2024b), since a number of laws require universal opt-out options.
@@
To make it applicable to various laws the setting would need to be described to users in a wording that is not specific to one particular law, e.g., “you can enable the toggle to exercise your opt-out right to the extent of the law applicable in your region.” The description could also be custom-tailored to specific laws if it is known which laws are applicable from account settings or location detection mechanisms. Before any selling or sharing of personal information could occur, third-party libraries could query a user’s opt-out status via a call to the new opt-out right API, as re-purposed from the current AdID API. If not yet known at the time of the call, a user should be prompted to select a preference. If users opt out, they should be opted out to the maximum extent per the applicable law.
@@
Exercising the opt-out right may have impact on other platform-level privacy settings, which would need to be aligned. Notably, while its further direction is not yet fully clear (Chavez, 2025), Google’s Privacy Sandbox, a privacy framework for the web and Android, includes various settings that could be impacted. In general, if users exercise their opt-out right, any setting that would run contrary to what the right requires would need to be turned off, e.g., under the CCPA, settings that would enable the selling or sharing of personal information. The opt-out right setting should also be available to users on devices without Google Play Services, i.e., on alternative Android distributions. As the simplest solution, those could re-implement the Google APIs as, e.g., GrapheneOS (GrapheneOS, 2024) and /e/OS (e Foundation, 2024), already do. Overall, the emerging legal landscape requiring online ads to be served in a privacy-preserving manner presents an opportunity for Google to re-purpose the AdID setting towards the opt-out right and align them with the mandatory legal requirements in California and other jurisdictions.
@@ 7.4. Legislative and Regulatory Support
In our view, the platform-level implementation of an opt-out right would benefit from additional legislative and regulatory support. The online ad industry can be resistant to change. A frictionless and universal opt-out setting presented prominently to users would likely see a high adoption rate as was the case for Apple’s App Tracking Transparency framework (Kollnig et al., 2022b). Such setting would reduce the amount of data available for ad personalization and potentially impact ad networks’ revenue given their current business model. Historically, the online ad industry has resisted changes that limit data access.
@@
On the legislative side a promising proposal had been adopted in California with Assembly Bill 3048 (California Legislature, 2024). This Bill required Google and other companies to implement in mobile operating systems a setting that enables consumers to send opt-out preference signals. However, following lobbying by the online ad industry, the California Governor chose to veto the Bill stating that “to ensure the ongoing usability of mobile devices, it’s best if design questions are first addressed by developers, rather than by regulators” (Office of the Governor, 2024). Platform operators have, however, not taken on the task, and we are not aware of major efforts currently under way.
@@
After Assembly Bill 3048 failed a new Assembly Bill 566 was introduced (California Legislature, 2025). This Bill is now pending as the California Opt Me Out Act. It requires browser vendors to “include a setting that enables a consumer to send an opt-out preference signal to businesses with which the consumer interacts through the browser.” Notably, the language of the former Assembly Bill 3048 on mobile app operating systems is absent. [^29] The current text of Assembly Bill 566 also covers “browser engines,” which is intended to include WebViews and similar web technologies on mobile devices. However, this wording will be removed going forward (Bender, 2025). Thus, even if the California Opt Me Out Act becomes law, it will not improve the situation on mobile platforms.
@@
However, even without further legislative action, regulatory agencies have an opportunity to step in. Notably, the California Privacy Protection Agency (CPPA) has published the CCPA Regulations that state that consumers must be able to opt out of the sale and sharing of their personal information without having to make individualized requests with each business. [^30] CCPA Regulations, §7025(a). Consumers have a right to opt out on every “device,” i.e., “any physical object that is capable of connecting to the internet, directly or indirectly, or to another device.” [^31] CCPA, §1798.140(o). To achieve this goal the CPPA can “[p]rovide guidance to businesses regarding their duties and responsibilities.” [^32] CCPA, §1798.199.40(f). Thus, the CCPA Regulations provide the CPPA with a tool to guide platform operators toward the implementation of an opt-out right setting. We think that this regulatory support would help to establish the CCPA opt-out right on Android and other platforms.
@@ 8. Conclusions
We study the effectiveness of opting out per the CCPA from ad tracking via GPC and app-level opt-out UIs. We find that the CCPA opt-out right applies to 62%–81% of the top-free Android apps on the US Play Store and that many apps are not compliant. For example, when sending GPC signals and disabling apps’ access to the AdID, 338 apps still had the ccpa status of the ad network Vungle set to opted in while only 26 had set it to opted out. The main problem is that the opt-out right is not supported by a platform-level setting. Moving forward, as the AdID setting is not aligned with users’ opt-out right, the current AdID setting should be re-purposed towards an opt-out right setting. As we focused on Android, the opt-out compliance on iOS and other platforms remains an open question. It may be that the App Tracking Transparency framework (Apple Developer, 2024) does not help users with their opt-out right either resulting in similar compliance rates as we found here for Android. In general, it would be a worthwhile endeavor to evaluate the extent to which opt-out rights are respected on different platforms and for different laws.
9. Ethical Principles
We designed and conducted our study to minimize risk and adhere to ethical research principles. All data collection was limited to the personal information of the researchers. As the study did not involve external human subjects, it did not require review by our institutions’ Institutional Review Board. The impact on the evaluated apps’ infrastructure was minimal as it was limited to approximately 30 seconds per app. In line with responsible disclosure principles, we have notified all companies mentioned by name in this paper whose services were observed accessing the Android AdID after it had been disabled. We are also in the process of contacting app operators who stated that the CCPA opt-out right is available to California residents in their privacy policy but failed to provide a corresponding setting. The use of Mitmproxy for the purpose of our research is a standard and common methodology (Kollnig et al., 2022a).
Acknowledgment
We thank the anonymous reviewers for their valuable suggestions. Don Marti and Aram Zucker-Scharff provided excellent feedback and industry expertise on our draft. We are grateful to the National Science Foundation (Award #2055196) and to the Alfred P. Sloan Foundation (Program in Universal Access to Knowledge) for their support of this research. We thank Wesleyan University, its Department of Mathematics and Computer Science, and the Anil Fernando Endowment for their additional support. We also thank Wesley Tan for support of the research in its earlier stages. Conclusions reached or positions taken are our own and not necessarily those of our financial supporters, its trustees, officers, or staff.

### v6

@@ 6.2.2. Connecting to Ad Tracking Domains
Figure 9. Counts of apps in our app dataset connecting to the 29 most prevalent ad tracking domains that are subject to the CCPA opt-out right (Figure 5). There are no substantial differences across conditions.:
Evaluating the counts of apps in our app dataset that make ad tracking domain connections, we do not find substantial differences across conditions. However, there are minor differences. Notably, the count of apps that connect to the app-measurement.com domain in the AdID condition is higher (1,047 apps) than in the No AdID + GPC condition (808 apps). The reason for this difference could be a compounding effect where disabling the AdID and sending GPC signals each leads to a slight decrease. Overall, however, the number of apps connecting to ad tracking domains in the different conditions does not indicate that sending GPC signals is an effective means for exercising the CCPA opt-out right. The same is true for disabling the AdID. Figure 9 illustrates our results.
One other point is noteworthy. In our app dataset, all apps that connected to first-party servers also connected to third-party servers. However, 12% of apps connected exclusively to third-party servers. Those apps will receive GPC headers directly from a user’s device without reliance on redirects from a first-party server. For those apps we did not observe substantially different levels of GPC compliance. Thus, it does not appear that the communication model—receiving GPC signals via redirects from first parties vs direct GPC signals from a user’s device—plays a role for GPC compliance. This is in line with the GPC specification, which does not differentiate between the two (Zimmeck et al., 2025). Third parties are obligated to listen for GPC signals regardless from whom they come.
@@ 6.2.3. Disclosing Device Identifiers
Overall, the rates of device identifier disclosures under the four conditions are similar. The Android ID, serial number, and MAC address of our test device was not disclosed in any condition. In some requests to googleapis.com, we found our test device’s IMEI, which, however, is not correlated to any particular condition. Further, we found the public IP address of our test device in some third-party request or response bodies, though, again, not correlated to any particular condition. Interestingly, a small number of domains still received the AdID in the No AdID and No AdID + GPC conditions. [^27] Three of the domains are Google domains — app-measurement.com, googleapis.com, and google-analytics.com — and three are non-Google domains — omtrdc.net, gadsme.com, and wizz.chat. We re-ran our analysis, which confirmed our results, leaving us uncertain about the reason.
Since we quit each app and cleared its cache and data in between conditions, the AdID from a previous condition run could not have been left stored in the app’s storage at the time of the next condition run. Generally, the app could also not have written the AdID to other apps’ external storage on our test device. While our controller granted each app all permissions it requested at install time, the app would generally need direct user consent to read and write to other apps’ files (Android Open Source Project, 2025b). A plausible explanation is that the AdID was stored at the backend and associated with another identifier that was then synced once the app was run again in the new condition. We inquired with the respective companies but did not hear back.
@@ 7. Discussion: Implementing Opt-out Rights on Android and Improving User Privacy
Our results suggest that many Android apps do not provide California residents with an effective way to exercise their CCPA opt-out right. We see this shortcoming as a platform-level challenge. To broadly improve the privacy of Android users the Android platform would benefit from an opt-out setting to exercise opt-out rights under applicable law in place of its current AdID setting.
@@ 7.1. Android’s Opt-out Right Gap
As we have seen (Section 6.2), the long-established Android AdID setting has no substantial effect on apps’ opt-out behavior. Apps’ sharing and selling of personal information also generally continues despite sending GPC signals. At the app-level, our results suggest that many apps do not have a dedicated CCPA opt-out setting, and some that do are not propagating California residents’ opt-out preferences to third parties (Section 5.2). Given that 63%–82% of apps in our app dataset are required to respect the CCPA opt-out right (Section 6.1), our results suggest that many apps violate the CCPA as they do not enable California residents to exercise their right. Our results suggest that many apps do not provide functionality for California residents to exercise their opt-out right.
There could be various reasons for why that is the case. First, traditionally the core business model of the ad online ecosystem has been directly opposed to respecting opt-out preferences (Zuboff, 2019). Until the recent enforcement actions by the Office of the California Attorney General and other attorneys general offices (State of California Department of Justice, 2022b, a, 2025) the risk of being caught for not doing so was low. Some app operators may simply accept the risk. After all, dark patterns are ostensibly proliferating because companies’ A-B testing has revealed them to be profit maximizing (Luguri and Strahilevitz, 2021). The same could be true for not providing opt-out functionality.
In our view, technical and ecosystem complexities also play a major role for the lack of opt-out right compliance. Opting out is a complex coordination problem with distributed responsibilities among first parties, ad networks, CMPs, and other parties, all of whose activities and motivations are not aligned (Hils et al., 2021). These different responsibilities become difficult to manage without addressing them at the platform-level. Thus, in our view, our results indicate that app-level opt-out preferences are just not feasible. We need a platform-level opt-out right architecture.
@@ 7.2. Two Settings for One Task
As it stands, Android has two types of opt-out settings: the platform-level AdID and app-level CCPA UI opt-out settings. The AdID setting is not intended for exercising the opt-out right per the CCPA or other privacy laws. However, according to Google, if a user has disabled the AdID, operators “may not use the advertising identifier for creating user profiles for advertising purposes or for targeting users with personalized advertising” (Google, 2025e). Similarly, the CCPA prohibits the sharing of personal information for purposes of cross-context behavioral advertising, which is defined as the “targeting of advertising to a consumer based on the consumer’s personal information” and which covers “consumer profiles” per the CCPA Regulations. [^28] CCPA §1798.140(k), CCPA Regulations §7025(c)(1). Thus, the intended effect of the AdID setting is similar to what the CCPA opt-out right is intended for. However, despite this similarity the AdID setting is without legal meaning outside the contractual relationship that Google has with the app operators.
This separation of the AdID setting from the opt-out right is based on its historical development predating the introduction of the CCPA (Section 2.2). In our view, it would benefit the Android platform if both were united and the opt-out right would be available on Android as a native setting. There should be only one opt-out setting on Android. The current separation is confusing from a user perspective. Users have to both disable the AdID and perform app-level opt-outs to accomplish one task. If there were only one setting, it would give users a consistent experience and also relieve app operators from compliance burdens they may not be sufficiently equipped to tackle. Thus, we propose to re-purpose the existing Android AdID setting as an opt-out right setting with legal meaning under the CCPA and other privacy laws.
@@ 7.3. Re-purposing the AdID Setting
Disabling the AdID should be interpreted as exercising the opt-out right. This would immediately make the opt-out right on Android available to a wide audience in line with previous results suggesting that many users have interest in such setting (Zimmeck et al., 2023). Such re-purposing would unify and streamline the current two types of settings—the platform-level AdID setting and the app-level opt-out right settings—into one and lead to greater legal clarity. The current AdID API can be modified to return an opt-out signal (akin to GPC’s Sec-GPC: 1 header) instead of a string of zeros when a user opts out (Google, 2025a). Notably, opting out via GPC and the AdID setting are both intended as universal settings and have nearly identical scope (Appendix 11.2). However, ultimately, whether this uniform setting is a GPC setting or other opt-out right setting is secondary.
@@
Evolving the AdID setting and APIs towards the opt-out right requires a combination of technological and policy changes on the Android platform. However, creating a platform-level API that apps and their integrated libraries can query for the user’s opt-out status is a standard feature for an operating system or other platform-level software layer. To make users aware of their opt-out right, Android should surface a prompt for universally opting out as part of the setup of a new or reset Android device. Android could also surface such prompt upon the first start of an app that is selling or sharing personal information. Crucially, such prompt should contain an option for applying a user’s preference universally across all apps, e.g., via generalizable active privacy choice (Zimmeck et al., 2024b), since a number of laws require universal opt-out options.
@@
To make it applicable to various laws the setting should be described to users in a wording that is not specific to one particular law but broadly covers all potentially applicable laws, for example, “you can enable the toggle to exercise your opt-out right to the extent of the law applicable in your region.” The description could also be custom-tailored to specific laws if it is known which laws are applicable from a user’s account settings, via the user’s location, or otherwise. Before any selling or sharing of personal information could occur, third-party ad networks need to query a user’s opt-out status via a call to the new opt-out right API, as re-purposed from the current AdID API. If not yet known at the time of the call, a user should be prompted to select a preference, again, with a universal option. If users opt out, they should be opted out to the maximum extent per the applicable law.
@@
Exercising the opt-out right may have impact on other platform-level privacy settings, which would need to be aligned. Notably, while recently pared-back substantially (Google, 2025g), Google’s Privacy Sandbox, a privacy framework for the web and Android, includes various settings that could be impacted. In general, if users exercise their opt-out right, any setting that would run contrary to what the right requires would need to be turned off, e.g., under the CCPA, settings that would enable the selling or sharing of personal information. The opt-out right setting should also be available to users on devices without Google Play Services, i.e., on alternative Android distributions. As the simplest solution, those could re-implement the Google APIs as, e.g., GrapheneOS (GrapheneOS, 2025) and /e/OS (e Foundation, 2024), already do. Overall, the emerging legal landscape requiring online ads to be served in a privacy-preserving manner presents an opportunity for Google to re-purpose the AdID setting towards the opt-out right and align it with the mandatory legal requirements in California and other jurisdictions.
@@ 7.4. Legislative and Regulatory Support
In our view, the platform-level implementation of an opt-out setting would benefit from additional legislative and regulatory support. The online ad industry has shown to be resistant to change in the past. A frictionless and universal opt-out setting presented prominently to users would likely see a high adoption rate as was the case for Apple’s App Tracking Transparency framework (Kollnig et al., 2022b). Such setting would reduce the amount of data available for ad personalization and potentially impact ad networks’ revenue given their current business model.
@@
On the legislative side a promising proposal had been adopted in California with Assembly Bill 3048 (California Legislative Information, 2024). This Bill required Google and other companies to implement in mobile operating systems a setting that enables consumers to send opt-out preference signals. However, following lobbying by the online ad industry, the California Governor chose to veto the Bill stating that “to ensure the ongoing usability of mobile devices, it’s best if design questions are first addressed by developers, rather than by regulators” (Office of the Governor, 2024). Platform operators have, however, not taken on the task, and we are not aware of major efforts currently under way.
@@
After Assembly Bill 3048 failed, a new Assembly Bill 566, the California Opt Me Out Act, was introduced (California Legislative Information, 2025). This time the Bill was signed by the California Governor and will become law in January 2027 amending the CCPA. It requires businesses to “not develop or maintain a browser that does not include functionality configurable by a consumer that enables the browser to send an opt-out preference signal.” However, the language of the former Assembly Bill 3048 on mobile app operating systems is absent. [^29] A former version of Assembly Bill 566 also covered “browser engines,” which was intended to include WebViews and similar web technologies on mobile devices. However, this wording was removed from the final version (Bender, 2025). Thus, the California Opt Me Out Act will generally not improve the situation for users of mobile platforms.
@@
However, even without further legislative action, regulatory agencies have an opportunity to step in. Notably, the California Privacy Protection Agency (CPPA) has published the CCPA Regulations stating that consumers must be able to opt out of the sale and sharing of their personal information without having to make individualized requests with each business. [^30] CCPA Regulations §7025(a). Consumers have a right to opt out on every “device,” i.e., “any physical object that is capable of connecting to the internet, directly or indirectly, or to another device.” [^31] CCPA §1798.140(o). To achieve this goal the CPPA can “[p]rovide guidance to businesses regarding their duties and responsibilities.” [^32] CCPA §1798.199.40(f). Thus, the CCPA Regulations provide the CPPA with a tool to guide platform operators toward the implementation of an opt-out right setting. We firmly believe that this regulatory support would help to establish the CCPA opt-out right on Android and other platforms.
@@ 8. Conclusions
We studied the effectiveness of opting out per the CCPA from ad tracking via GPC and app-level opt-out UIs. We find that the CCPA opt-out right applies to 62%–81% of the top-free Android apps on the US Google Play store and that many apps are not compliant. For example, when sending GPC signals and disabling apps’ access to the AdID, 338 apps still had the ccpa status of the ad network Vungle set to opted_in while only 26 had set it to opted_out. The fundamental challenge is that the opt-out right is not supported by a platform-level setting. Moving forward, as the current AdID setting is not aligned with users’ opt-out right, it should be re-purposed towards an opt-out right setting. As we focused on Android, the opt-out compliance on iOS and other platforms remains an open question. It may be that the App Tracking Transparency framework (Apple Developer, 2025) does not help users with their opt-out right either resulting in similar levels of non-compliance as we found here for Android. In general, it would be a worthwhile endeavor to evaluate the extent to which opt-out rights are respected on different platforms and under different laws as well as to measure the impact of legislative and regulatory activities.
9. Ethical Considerations
We designed and conducted our study to minimize risk and adhere to ethical research principles. All data collection was limited to the personal information of the researchers. As the study did not involve external human subjects, it did not require Institutional Review Board approval. The impact on the evaluated apps’ infrastructure was minimal as it was limited to approximately 30 seconds for the initialization and each test condition per app. In line with responsible disclosure principles, we have notified all companies whose services we observed accessing the Android AdID after we had disabled it. We also contacted the app operators we identified in our UI opt-out evaluation who stated that the CCPA opt-out right is available to California residents in their privacy policy but failed to provide a corresponding setting in their app.
10. Availability of Artifacts
Setup instructions and scripts for performing the described dynamic analysis are publicly available at https://github.com/privacy-tech-lab/gpc-android.
Acknowledgements.
We thank our anonymous reviewers and our shepherd for their improvement suggestions. Don Marti and Aram Zucker-Scharff provided valuable feedback and industry expertise on our draft. We also thank former Wesleyan University students Wesley Tan, Samir Cerrato, Eliza Kuller, and Bella Tassone for support of the research in its earlier stages. We are grateful to the National Science Foundation for their support of this research (Award #2055196). We also thank Wesleyan University, its Department of Mathematics and Computer Science, and the Anil Fernando Endowment for their additional support. Conclusions reached or positions taken are our own and not necessarily those of our supporters, its trustees, officers, or staff.